# Does CLIP perceive art the same way we do?

### v1

@@ 4.3. Semantic Relationships in the Latent Space
Figure 5. Three-dimensional projection of the textual embeddings of artistic styles and the visual embeddings of correctly classified (green) and misclassified (red) images using UMAP.
To gain a deeper understanding of the structure of CLIP’s latent space, we conducted a detailed analysis based on nearest neighbors. Specifically, we focused on image pairs in which one image was correctly classified while the other was not. By identifying such pairs, we were able to isolate semantically coherent examples where classification performance diverged. This setup offered valuable insight into how subject matter similarity and stylistic cues interact within CLIP’s representation space, and where the model may struggle to disentangle the two.
Representative examples are shown in Figure 6, where the misclassified image is displayed on the left, and its nearest correctly classified neighbor appears on the right. For each image, we report the cosine similarity scores with respect to both the true and the predicted style prompts, computed in both the original latent space and its lower-dimensional projection. This comparison allows us to examine how proximity in the embedding space relates to classification outcomes and whether stylistic distinctions are preserved through dimensionality reduction.
Figure 6. Visual comparison between a misclassified image (left) and a correctly classified image (right). For each artwork, the actual and predicted artistic styles are shown, along with their respective similarity scores to the image in the latent CLIP space.
These findings underscore a key limitation of CLIP: it tends to encode and prioritize semantic content—like objects, scenes, and compositions—over stylistic features such as brushwork, color palette, or compositional structure. This leads to frequent misclassifications when artworks differ in style but share similar subject matter. This bias is consistent with CLIP’s training objectives, which favor semantic alignment based on large-scale image-text data, where captions often emphasize content over formal style. As a result, CLIP’s latent space tends to conflate stylistically diverse images with similar semantics. While we explored projections to mitigate this issue, the results were unsatisfactory. Techniques like light adapters could potentially fine-tune CLIP’s vision encoder; such modifications fall outside the scope of our current research aims.
5. Experiments with AI-generated artworks
@@
The same ambiguity arises in style classification: in the AI-Pastiche dataset, the ”style” corresponds to the intended style described in the prompt—not necessarily the one successfully rendered in the generated image. As such, any misclassification could be due either to the image generator failing to follow the prompt or to CLIP failing to recognize the intended style.
To clarify the contribution of these factors, we further compared CLIP’s perception of generated images with human judgments, drawing on user survey data from the AI-Pastiche dataset (Asperti et al., 2025a). Section 6 provides details of these experiments.
5.1. Image-prompt similarity
Here, we are comparing the embedding of the generated image with the embedding of the relative prompt. AI-pastiche prompts are quite articulated, and we faced a similar problem to the case of real image descriptions due to the maximum numbers of tokens accepted by CLIP. For this reason, we created a prompt-summary in a way similar to the technique described in Section 4.1.
We generated embeddings for both images and prompt summaries and computed their cosine similarity, using this value to predict the prompt for each generated image. Due to the smaller number of textual prompts (72), we just measure accuracy in this case.
The results are shown in Table 3.
@@ 5.1. Image-prompt similarity
Table 3. Accuracy of different CLIP models in matching generated images with their corresponding summarized prompts in the AI-Pastiche dataset.
@@
Table 4. Result table for art style recognition on AI-Pastiche
@@
All CLIP models perform well in associating generated images with their corresponding prompt summaries, with accuracy values exceeding 0.86 across the board. The highest performance is achieved by RN50x64, ViT-L/14, and ViT-L/14@336px, all reaching an accuracy of 0.896. The task is sensibly simpler than the image-description alignment of Section 4.1, since we only have 73 prompts relative to quite different subjects. Nevertheless, the results confirm CLIP’s ability to capture the visual-semantic correspondence in synthetically generated image-text pairs.
@@
From the perspective of the generators, the high recall indicates that all models performed well in producing images that closely matched the subjects described in the prompts. The average cosine similarity between each generated image and its corresponding prompt summary is 0.278, with a standard deviation of 0.344.
@@
Although cosine similarity can be profitably used to associate an image to its prompt, it is not clear if it can be reliably employed as a standalone metric to assess the quality of different generated images in terms of their alignment to the same prompt. The problem is that this assessment requires a complex evaluation comprising not just the semantic correspondence with the subject but also the stylistic adherence and the technical quality of the generation. This includes evaluating the absence of artifacts, distortions, or visual defects that may not be compatible with the intended artistic style.
@@
We start addressing stylistic issues in Section 5.2; and in Section 6.1 we will compare the CLIP-evaluation of the adherence between an image and its prompt with a similar evaluation done by human experts.
5.2. Style Recognition
@@
In the second experiment, we used CLIP to evaluate the alignment between generated artworks and their expected styles, provided among the AI-Pastiche metadata. Similar to the case of images from NGAD, we generated a prompt of the form ”an artwork in [style] style” and computed its cosine similarity with the image embedding, deriving a confusion matrix. The accuracy results for the different models are shown in Table 2.
Considering that we are comparing the predicted style of generated images with the target style specified in the prompt, the classification accuracy achieved by CLIP is surprisingly high and comparable to the accuracy observed in the NGAD task.
Upon visual inspection, the generators evaluated in the AI-Pastiche dataset do not appear to reproduce the historical styles specified in the prompts convincingly.
In this context, the high agreement between CLIP’s classification and the original prompt may reflect a shared inductive bias rather than genuine stylistic fidelity: the reason is that many generative models rely on CLIP during training or inference as a scoring function, conditioning mechanism, or similarity guide.
In analogy with the work done for NGAD, a confusion matrix was constructed to better analyze CLIP model for predicting style in AI-Pastiche (Figure 7) and the distribution histograms of the actual and predicted values for AI-Pastiche (Figure 8).
Figure 7. Confusion matrix of the best CLIP model for predicting style in AI-Pastiche
Figure 8. Distribution of actual values and predicted values in AI-Pastiche
@@
To better understand how CLIP assessments compare with human perception, we turn to the adherence and artifact surveys provided in the AI-Pastiche dataset. These human evaluations offer a valuable reference point for gauging whether CLIP captures stylistic and visual cues in a way that aligns with human judgment.
6. Comparison with human evaluations
In this section, we directly compare CLIP’s perceptual judgments with human interpretations, drawing on the results of two user surveys from the AI-Pastiche dataset: the adherence survey, which measures how closely generated images align with their prompt, and the artifact survey, which evaluates the presence of visual distortions or inconsistencies. These comparisons provide a human-centered benchmark for evaluating CLIP’s ability to detect stylistic cues and perceptual artifacts in generated artwork.
6.1. Adherence Analysis
@@
In (Asperti et al., 2025a), an adherence survey was conducted in which participants were shown a set of images generated by different models from the same textual prompt. They were asked to rate each image as good, neutral, or bad in terms of its adherence to the prompt in a comparative setting. By averaging responses from multiple participants, the authors derived an adherence score for each image, reflecting perceived stylistic and semantic alignment with the original prompt.
@@
To explore whether CLIP captures similar perceptual cues, we conducted a parallel analysis using CLIP similarity scores. For each prompt, we computed the cosine similarity between the CLIP text embedding of the prompt and the CLIP image embeddings of all associated generated images. These scores were then normalized and scaled to produce a vector directly comparable to the human-derived adherence scores.
@@
Finally, we computed the correlation between the CLIP similarity vector and the human adherence scores, using it as a synthetic measure of alignment between CLIP’s assessments and human perception. This procedure was repeated across all available CLIP models, and the results are reported in the second column of Table 5 (the third column is discussed in Section 6.2).
Model Correlation with human evaluation Integration of defects
RN50 0.406 0.478
RN101 0.398 0.462
RN50x4 0.379 0.448
RN50x16 0.413 0.489
RN50x64 0.428 0.484
ViT-B/32 0.411 0.481
ViT-B/16 0.383 0.458
ViT-L/14 0.425 0.482
ViT-L/14@336px 0.437 0.497
@@
Table 5. Correlation between a human evaluation of the adherence between a generated image and its prompt, compared with a similar evaluation based on CLIP’s.
Across all models, the similarity scores are around 0.4. The result is not too low, but it should be compared with the correlation between evaluations of different humans, which on average is around 0.7, which is sensibly higher.
6.2. Perception of Artifacts and Deformations
According to the instruction provided in (Asperti et al., 2025a), the user was asked to perform an adherence evaluation essentially driven by three different directions:
  1. (1)
    Content: respect of the subject described in the prompt:
  2. (2)
    Style: respect of the style that the generative model was required to imitate;
  3. (3)
    Overall Quality: absence of visible artifacts and deformations typical of generative techniques.
@@ 6.2. Perception of Artifacts and Deformations
One of the surveys collected in (Asperti et al., 2025a) was specifically aimed at detecting the presence of visible artifacts or deformations in the generated image. Defects were categorized as Major (clearly visible of frequent errors, such as macroscopic anatomical mistakes), Minor (additional fingers, minor deformations), or None (no apparent mistake). Results were summarized in a ”defect score” associated with each image.
@@
We sought to investigate whether CLIP could identify and detect these kinds of mistakes. Our initial investigation aimed to determine if we could approximate the defect score through linear regression, starting from the CLIP embedding of the AI-Pastiche images.
The result was negative: we obtained a coefficient of determination $R^{2}$ close to 0.
@@
As additional evidence that CLIP embeddings do not account for defects and artifacts in input images, we test whether CLIP’s evaluation of prompt adherence can be improved by incorporating a linear combination with the human-evaluated ”defect score”. This turns out to be the case: a suitable linear combination achieves a similarity of nearly 0.5 with the average human adherence evaluation (see the third column in Table 5).
Nevertheless, CLIP sometimes assigns high adherence scores to images that humans evaluate poorly and vice versa. This discrepancy is often due to mismatches in content or style. Some illustrative examples are shown in Figure 9.
@@
Figure 9. Examples of images in the AI AI-Pastiche dataset not aligning well with their prompts, for content or style.
Figure (a) was supposed to represent a fight between two knights on horseback; humans penalized the fact that only one knight is represented, while CLIP does not seem to notice the absence. Image (b), generated by Auto-Aesthetics V1, was meant to depict a lively Parisian street scene in the rain, in a style resembling impressionism: humans reproach the lack of adherence to the required style. The prompt for image (c) required the representation of a kneeling figure in rich robes offering a white flower to another figure, draped in blue; the fact that one of the two figures is missing is not captured by CLIP similarity.
7. Conclusion
@@
Our investigation into CLIP’s perception of artworks—across both human-created and AI-generated images—reveals a model with remarkable breadth, but one still far from capturing the richness of human aesthetic and contextual understanding. While CLIP is adept at grounding images in broad semantic categories and descriptive summaries, it often falters when asked to navigate the more subjective terrain of artistic nuance: style, intention, emotional tone, and cultural context.
This gap is particularly evident in generative contexts, where CLIP must assess not only what is depicted, but how and why. When the prompt serves as both the generative instruction and the evaluative anchor, the act of measuring alignment becomes ambiguous: are mismatches due to the generator’s limitations or CLIP’s perceptual blind spots? Our experiments, informed by human judgments, suggest that CLIP often lacks the sensitivity required to discern these subtle failures—especially when artifacts are stylistically ”plausible” but semantically off-key.
More fundamentally, our work raises questions about how we interrogate multimodal representations. CLIP’s joint embedding space promises a bridge between vision and language, yet that bridge is asymmetrical: text embeddings are inherently structured and interpretable, while image embeddings remain opaque and spatially abstract. Probing image understanding through text similarity metrics can therefore be misleading, especially when the image conveys signals that elude linguistic capture—such as texture, composition, or emotional affect. In the case of art, where meaning is often non-verbal and deliberately ambiguous, this mismatch becomes a critical limitation.
@@
Looking ahead, we believe that future vision-language systems must go beyond mere alignment. What is needed is a deeper model of perception—one that can reason about images not just in terms of objects or styles, but in terms of historical context, artistic intent, and visual storytelling. This may require models trained with richer supervision, involving not only captioned data but also art historical metadata, expert narratives, and multimodal dialogues.
Ultimately, as AI is increasingly used to create, curate, and critique visual culture, we must ask not just what models see but how they see, and whose eyes they are borrowing. CLIP represents a powerful beginning, but it is not yet a substitute for human perception in the arts. Rather, it is a lens—partial, biased, but illuminating—that can help us better understand both machine vision and our own ways of seeing.
Acknowledgements.
This Research has been partially supported by the Future AI Research (FAIR) project of the National Recovery and Resilience Plan (NRRP), Mission 4 Component 2 Investment 1.3 funded from the European Union - NextGenerationEU.
8. Appendices
Appendix A Example of an artwork from the National Gallery
To illustrate the methodology adopted in our study, we selected a painting from the National Gallery of Art, Van Gogh’s self-portrait. Since zero-shot CLIP accepts textual descriptions of no more than 77 tokens, we employed GPT-4o mini to generate a summary of the description of the painting, restricted to a maximum of 300 characters. The summarization process aimed to preserve essential information related to the subject and the stylistic characteristics of the artwork. This is the actual prompt used to generate the summaries:
Your goal is to summarize the following painting descriptions in 300 characters.
You will be provided the description of a painting, its subject, its style, and its period, and you will output a JSON object containing the following information:
summary: string // at most 300 characters summary of the painting based on the painting description.
The summary must retain information about the subject, style, and period.
In Figure  10 we present the painting, the official description provided by the National Gallery, and the summarized version produced using GPT-4o mini.
Figure 10. Example of an artwork from the National Gallery.
National Gallery Description:
Shown from the chest up, a man with short, orange hair and green-tinted, pale skin looks at us, wearing a vivid blue painter’s smock in this vertical portrait painting. His smock and the background are painted with long, mostly parallel strokes of cobalt, azure, and lapis blue. His shoulders are angled to our left, and he looks at us from the corners of his blue eyes. He has a long, slightly bumped nose, and his lips are closed within a full, rust-orange beard. He holds a palette and paintbrushes in his left hand, in the lower left corner of the canvas. The background is painted with long brushstrokes that follow the contours of his head and torso to create an aura-like effect.
Summarized version of the description:
This Post-Impressionist portrait by Vincent van Gogh (1876-1900) depicts a man with orange hair and pale skin in a blue painter’s smock. He gazes at the viewer with blue eyes, holding a palette and brushes. The background features long, parallel strokes of blue, enhancing the aura around him.
Appendix B Example of an AI-generated artwork from the AI-Pastiche Dataset
As an example of the AI-Pastiche dataset, we chose an Impressionist painting generated by generative models, depicting a bridge landscape. As observed in the artwork, the generative model adhered to the given prompt, producing an image strongly reminiscent of Monet’s Japanese bridge. For the CLIP processing, the same procedure was applied to the description, limiting it to 300 characters of text, as done for the National Gallery descriptions.
In Figure  11, we present the generated painting, the prompt provided to the model, and the corresponding summary.
Figure 11. Example of an artwork from the AI-Pastiche.
Prompt for the generated image:
Create an Impressionist-style painting depicting a serene outdoor scene, such as a sunlit garden, a riverside, or a city park. The image should focus on capturing the play of natural light and atmosphere, with soft, loose brushstrokes and a pastel-like color palette of light blues, greens, pinks, and yellows. The figures and landscape should appear slightly blurred, as if seen from a distance, giving a sense of movement and fleeting moments. Include reflections in water, dappled sunlight, and subtle shifts in color to evoke a peaceful, idyllic mood, typical of Impressionist art.
Summarized version of the prompt:
This Impressionist painting from the XIX century features a serene landscape with soft tones, depicting a sunlit garden or riverside. It captures natural light and atmosphere through loose brushstrokes and a pastel palette. Figures and scenery appear slightly blurred, evoking movement and tranquility, with reflections in water.
Appendix C The distribution of labels predicted by all CLIP models
Figure 12 presents the style prediction results for all CLIP models. As shown in the graph, most models exhibit challenges in accurately identifying Romantic, Art Nouveau, and Dadaist styles. Notably, many models overestimated the prevalence of Romantic artworks compared to their actual representation in the dataset. The underperformance in predicting Art Nouveau and Dadaist styles may stem from their visual similarities to avant-garde movements such as Futurism, Abstract art, Cubism, and Surrealism, which could lead to misclassifications. However, as observed in our analysis, the ViT 14@336px model demonstrates the strongest overall performance, achieving higher accuracy and robustness across style categories compared to other architectures.
Figure 12. Histogram of style prediction of all CLIP models on AI-Pastiche
Appendix D CLIP Image embedding visualization
Figure 13. 3D UMAP projection of image embeddings of National Gallery of Art Dataset extracted from the CLIP ViT-L/14 model.
Each point represents a painting from the National Gallery dataset, colored by artistic style. The visualization reveals the semantic organization of the artworks in the latent space, where similar styles tend to cluster together.
Appendix E Example of a wrong summary associated with an artwork of the National Gallery of Art
Figure 14. Example of a misclassified summary
In Figure 14, there is an example of misclassifications in the summary recognition of an artwork in the National Gallery Dataset of Art computed using the CLIP ViT-L/14 model.
Actual summary:
In this Baroque painting by Canaletto (1751-1775), sunlight illuminates a vibrant landscape featuring an arched stone ruin, a bridge, and a river. Scattered figures in colorful attire engage with nature, while buildings and churches rise on a distant hill. The scene is rich with detail, showcasing Venetian life and architecture.
Predicted summary:
In this Rococo painting by Jean Honoré Fragonard (1751-1775), a lively scene unfolds in a lush park where light-skinned figures enjoy leisure by a river. A couple in elegant attire sits nearby, while boys engage in playful horse-riding games. Tall trees frame the idyllic landscape, enhancing the theme of amusement.
Appendix F Example of artifact presence in AI-generated artwork from the AI-Pastiche dataset
Figure 15 displays an AI-generated painting depicting Madonna and Child. As is characteristic of synthetic artworks produced by current generative models, the image exhibits pronounced anatomical distortions, particularly in the rendering of hands and facial features. A key objective of our study is to investigate whether CLIP-based analysis can directly extract meaningful information about such artifacts from image embeddings, bypassing manual annotation. By probing the latent representations of these synthetic outputs, we aim to identify patterns associated with structural anomalies (e.g., incoherent geometries and asymmetries) and evaluate their utility for automated quality assessment or iterative refinement of generative models.
Figure 15. AI-generated artwork from the AI-Pastiche dataset depicting Madonna and Child, showcasing anatomical artifacts.

### v2

@@ IV-C Semantic Relationships in the Latent Space
Figure 4: Three-dimensional projection of the textual embeddings of artistic styles and the visual embeddings of correctly classified (green) and misclassified (red) images using UMAP.
To better understand the structure of CLIP’s latent space, we performed a nearest-neighbor analysis, focusing on image pairs where one image was correctly classified and the other was not. These pairs isolate semantically coherent cases with divergent classification outcomes, revealing how subject similarity and stylistic cues interact and where the model struggles to disentangle them.
Representative examples are shown in Figure 5, with the misclassified image on the left and its nearest correctly classified neighbor on the right. For each image, cosine similarity scores are reported for both the true and predicted style prompts, in the original latent space and its lower-dimensional projection. This comparison highlights how embedding-space proximity relates to classification outcomes and whether stylistic distinctions are preserved after dimensionality reduction.
Figure 5: Visual comparison between a misclassified image (left) and a correctly classified one (right). For each artwork, the actual and predicted artistic styles are shown, along with their respective similarity scores to the image in CLIP’s space.
These findings highlight a key limitation of CLIP: it prioritizes semantic content - objects, scenes, and compositions - over stylistic features such as brushwork or color palette. This bias, rooted in training objectives that favor content-based alignment, often leads to misclassifications when artworks share subject matter but differ in style, causing the latent space to conflate stylistically distinct images with similar semantics.
V Experiments with AI-generated artworks
@@
TABLE III: Accuracy of different CLIP models in matching generated images with their corresponding summarized prompts in the AI-Pastiche dataset.
@@
TABLE IV: Result table for art style recognition on AI-Pastiche
The same ambiguity arises in style classification: in the AI-Pastiche dataset, the “style” corresponds to the intended style described in the prompt—not necessarily the one successfully rendered in the generated image. As such, any misclassification could be due either to the image generator failing to follow the prompt or to CLIP failing to recognize the intended style.
To clarify the contribution of these factors, we further compared CLIP’s perception of generated images with human judgments, drawing on user survey data from the AI-Pastiche dataset [3]. Section VI provides details of these experiments.
V-A Image-prompt similarity
Here, we are comparing the embedding of the generated image with the embedding of the relative prompt.
Embeddings were generated for both images and prompt summaries, and their cosine similarity was calculated. This similarity score was then used to predict the prompt corresponding to each generated image.
@@ V-A Image-prompt similarity
Due to the smaller number of textual prompts (72), only accuracy was measured in this case. The results are shown in Table III. All CLIP models perform well in associating generated images with their corresponding prompt summaries, with accuracy values exceeding 0.86 across the board. The highest performance is achieved by RN50x64, ViT-L/14, and ViT-L/14@336px, all reaching an accuracy of 0.896. The task is sensibly simpler than the image-description alignment of Section IV-A, since we only have 73 prompts relative to quite different subjects. Nevertheless, the results confirm CLIP’s ability to capture the visual-semantic correspondence in synthetically generated image-text pairs.
@@
From the perspective of the generators, the high recall indicates that all models perform well in producing images that closely match the subjects described in the prompts. The average cosine similarity between each generated image and its corresponding prompt summary is 0.278, with a standard deviation of 0.344.
@@
Although cosine similarity can be profitably used to associate an image with its prompt, it is not clear if it can be reliably employed as a standalone metric to compare the quality of images obtained from different generators on the same prompt. The problem is that this assessment requires a complex evaluation comprising not just the semantic correspondence with the subject but also the stylistic adherence and the technical quality of the generation. This includes evaluating the absence of artifacts, distortions, or visual defects that may not be compatible with the intended artistic style.
@@
We start addressing stylistic issues in Section V-B, and in Section VI-A we will compare the CLIP-evaluation of the adherence between an image and its prompt with a similar evaluation done by humans.
V-B Style Recognition
@@
In the second experiment, we used CLIP to evaluate the alignment between generated artworks and their expected styles, provided among the AI-Pastiche metadata. Similar to the case of images from NGAD, we generated a prompt of the form ”an artwork in [style] style” and computed its cosine similarity with the image embedding. The accuracy results for the different models are shown in Table II.
@@
Figure 6: Distribution of actual vs predicted styles in AI-Pastiche
CLIP achieves a surprisingly high accuracy in predicting the target style of generated images, comparable to that in the NGAD task, despite the fact that, on visual inspection, AI-Pastiche generators often fail to convincingly reproduce the historical styles specified in the prompts. This high agreement may stem from a shared inductive bias, since many generative models integrate CLIP as a scoring function, conditioning mechanism, or similarity guide. To evaluate whether CLIP’s style assessments align with human perception, the adherence and artifact surveys from the AI-Pastiche dataset provide a useful point of comparison.
VI Comparison with human evaluations
This section compares CLIP’s perceptual judgments with human evaluations from two AI-Pastiche surveys: adherence, measuring alignment with prompts, and artifact, assessing visual distortions.
VI-A Adherence Analysis
@@
In [3], an adherence survey asked participants to rate images generated from the same prompt as good, neutral, or bad in terms of prompt adherence. Averaging these ratings produced an adherence score for each image, reflecting perceived stylistic and semantic alignment.
@@
To assess whether CLIP captures similar cues, cosine similarity was computed between the CLIP text embedding of each prompt and the CLIP image embeddings of its generated images. Scores were normalized and scaled to match the human-derived adherence scores.
@@
The correlation between these CLIP scores and human ratings was then calculated as a synthetic measure of alignment, repeated for all available CLIP models. Results appear in the second column of Table V (the third column is discussed in Section VI-B).
human vs human CLIP vs human CLIP + defects vs human
0.70 0.43 0.50
@@
TABLE V: Correlation between humans’ evaluation of the adherence between a generated image and its prompt, and a similar evaluation based on CLIP’s (ViT-L/14@336px)
The average correlation between human evaluations of different individuals is around 0.7, while the correlation between CLIP’s adherence assessment with the average human assessment is sensibly lower: 0.43.
VI-B Perception of Artifacts and Deformations
@@
One of the surveys collected in [3] aimed to detect the presence of visible artifacts or deformations in the generated image. Defects were categorized as Major (clearly visible or frequent errors, such as macroscopic anatomical mistakes), Minor (additional fingers, minor deformations), or None (no apparent mistake). Results were summarized in a ”defect score” associated with each image.
@@
We sought to investigate whether CLIP could identify and detect these kinds of mistakes. Our initial investigation aimed to determine if we could approximate the defect score through linear regression, starting from the CLIP embedding of the AI-Pastiche images. The result was negative: we obtained a coefficient of determination $R^{2}$ close to 0.
@@
As additional evidence that CLIP embeddings do not account for defects and artifacts in input images, we tested whether CLIP’s evaluation of prompt adherence could be improved by incorporating a linear combination with the human-evaluated ”defect score”. This turns out to be the case: a suitable linear combination achieves a similarity of nearly 0.5 with the average human adherence evaluation (see the third column in Table V).
The remaining mistakes in CLIP’s perception seem to be mainly related to counting problems or stylistic issues. Some illustrative examples are shown in Figure 7.
@@
Figure 7: Examples of images in the AI AI-Pastiche dataset not aligning well with their prompts, for content or style.
Figure (a) was intended to depict two knights fighting on horseback; humans penalized the absence of one knight, which CLIP overlooked. Image (b), from Auto-Aesthetics V1, was meant to show a rainy Parisian street in an impressionist style; humans criticized its stylistic inaccuracy. For image (c), the prompt specified a kneeling figure in rich robes offering a white flower to another figure in blue; CLIP similarity did not register the absence of one figure.
VII Conclusion
@@
Our investigation into CLIP’s perception of artworks, conducted across both human-created and AI-generated images, reveals a model with remarkable breadth, but still far from capturing the richness of human aesthetics and contextual understanding. While CLIP is adept at grounding images in broad semantic categories and descriptive summaries, it often falters when asked to navigate the more subjective terrain of artistic nuance: style, composition, and technical accuracy.
Both the vision system and the cosine similarity measure have limitations. The vision system often fails to detect the presence and location of artifacts or defects. At the same time, extracting rich stylistic information through textual embeddings and simple cosine similarity is a challenge. This underuse of the visual representation is particularly problematic in art, where images convey signals - composition, narrative, emotional tone - that demand a deep semantic connection to their textual description.
@@
Looking ahead, future vision–language systems must go beyond mere alignment. What is needed is a deeper model of perception, one that can reason about images in terms of objects, styles, historical context, artistic intent, and visual storytelling. Achieving this may require richer supervision, incorporating not only captioned data but also art-historical metadata, expert narratives, and multimodal dialogues.
As AI takes on a greater role in creating, curating, and critiquing visual culture, the question becomes not just what models see, but how they see, and whose eyes they are borrowing. Current vision mechanisms are useful but partial lenses, limited in capturing the depth and nuance of human artistic perception. This study exposes such constraints in CLIP, highlighting the need to examine whether more advanced multimodal systems, including LLM-based architectures, can overcome these limitations or simply reproduce them in subtler forms.